Exercise 17-5 (Part Level Submission) On January 1, 2017, Metock Company acquires $140,000 of Spiderman Products, Inc., 10% bonds at a price of $136,579. Interest is received on January 1 of each year, and the bonds mature on January 1, 2020. The investment will provide Metlock Company a 11% yield. The bonds are classified as held to maturity▼ (b) Prepare a 3-year schedule o interest revenue and bond discount amortization, app ying he e ective-interest me od Run ansı ers 0 decima places. e.g 2 Schedule of Interest Revenue and Bond Discount Amortization Effective-Interest Method Bond Purchased to Yield Interest Bond Discount Carrying Amount Revenue Amortization Cash Date Received of Bonds 1/1/20 SHOW LIST OF ACCOUNTS LINK TO TEXT
Answer #1

Answer

  • All working forms part of the answer
  • Requirement

Date

Cash received

Interest Revenue

Bond Discount Amortisation

Carrying amount of Bonds

01-01-2017

$            136,579

01-01-2018

$                                      14,000

$                15,024

$                 1,024

$            137,603

01-01-2019

$                                      14,000

$                15,136

$                 1,136

$            138,739

01-01-2020

$                                      14,000

$                15,261

$                 1,261

$            140,000

---Working

Date

Cash received

Interest Revenue

Bond Discount Amortisation

Carrying amount of Bonds

01-01-2017

136579

01-01-2018

=140000*10%

=136579*11%

=15024-14000

=136579+1024

01-01-2019

=140000*10%

=137603*11%

=15136-14000

=137603+1136

01-01-2020

=140000*10%

=138739*11%

=15261-14000

=138739+1261